OpenStudy (anonymous):

f(x) = cos(x - 4) - 8 How do you do this? please explain with answer.

OpenStudy (anonymous):

How do you do what? Write instructions given.

OpenStudy (anonymous):

find the inverse function of the function f

OpenStudy (anonymous):

Solve for x y-8=cos x - cos 4 cox x = y - 8 + cos 4 \[x =\cos^{-1} (y-8+\cos 4)\]

OpenStudy (anonymous):

x=cos(y-4) -8 cos(y-4) = (x+8) y-4 = cos^-1 ( x+8) y= 4 + cos^-1 (x+8)

OpenStudy (anonymous):

\[f^{-1}(x) = 4 + \cos^{-1} (x+8)\]

OpenStudy (anonymous):

^ thats correct